OpenStudy (anonymous):

Factor the perfect square trinomial completely: 16y^2+56y+49

OpenStudy (anonymous):

Since you are given that this is a perfect square trinomial, what is: square root of (16y^2) = square root of (49) =

OpenStudy (anonymous):

4y and 7

OpenStudy (anonymous):

I guess I am just getting all of the ways to factor and procedures jumbled lol I had thought that B had to equal the sum of the factors of 49... :/

OpenStudy (anonymous):

ok.... so your answer is \(\large (4y + 7)(4y + 7) = (4y + 7)^2 \)

OpenStudy (anonymous):

B is the sum of 4*7 + 4*7 or 56 = 4*7 + 4*7